OpenStudy (anonymous):

Determine the zeros of f(x) = x3 – 12x2 + 28x – 9.

OpenStudy (phi):

at least this one has a 9, which does not have a zillion factors

OpenStudy (anonymous):

yeah

OpenStudy (amoodarya):

1st step try +-1 ,+-3 ,+-9

OpenStudy (anonymous):

@phi I got up to x^2-3x=-1 after synthetic division... Now I'm lost

OpenStudy (phi):

x^2 -3x+1=0 that does not factor into nice numbers. do you know the quadratic formula ?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

I got (3+/-√5)/2

OpenStudy (anonymous):

@phi I think I'm wrong

OpenStudy (phi):

that is correct. If you are sure you have the correct equation (no typos) then those are the roots: 9, \(\frac{3+\sqrt{5}}{2} \), \(\frac{3-\sqrt{5}}{2} \) you can use a calculator and test them

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

How do you do it if it's a quartic equation and not a cubic?

OpenStudy (phi):

you can still use the rational root theorem. But in the real world, we would use a computer and "root finding algorithms" http://en.wikipedia.org/wiki/Root-finding_algorithm